Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
bersanetti, casanueva - 20:23 Tuesday 14 October 2025 (67939) Print this report
Lock acquisition recovery after B1 PD2 replacement

After the replacement of B1 PD2 done today, we checked and recovered the lock acquisition:

  • the B1p_56MHz demodulation phase for DARM did not need any tuning;
  • the B1p_QD2_56MHz demodulation phase needed to be tuned; we made an initial tuning before closing the loop, then a finer one once in loop. Then later another step was done, only on TY, which improved the line separation between I and Q and increased the UGF, reducing then the loop gain (Figures).

Now _V and _H differ of 0.15 rad; to be observed in the following acquisitions if this behaviour is consistent and if it helps in reducing the big excursion of the loop gain at the beginning of CARM_NULL_1F that was observed yesterday, potentially causing issues.

We also removed the DARM offset reduction in ACQUIRE_LOW_NOISE_3 introduced yesterday, but it was left commented at line 6188 of ITF_LOCK.py in case.

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×